OpenStudy (anonymous):

Which is a solution to the equation 5x – y = –3?

OpenStudy (anonymous):

-21

OpenStudy (anonymous):

um. A. (2, –7) B. (1, 2) C. (–2, 7) D. (–2, –7)

OpenStudy (neer2890):

(-2,-7) put x=-2 and y=-7 in your equation you will get 5*(-2)-(-7)=-10+7= -3